# Disentangling Content from Style to Overcome Shortcut Learning: A Hybrid Generative-Discriminative Learning Framework

### v2

@@ 2.2.2 Identifying the Content Direction via Self-Distillation
While a conventional contrastive approach using an InfoNCE loss (Chen et al., 2020a) might seem applicable, we found it suboptimal for the fine-grained semantic alignment required between our strong positive pairs $(A,A^{\prime})$. This led us to reframe the problem: the core task is not to distinguish the pair from negatives, but to ensure their representations are identical in the content dimension.This objective aligns perfectly with the philosophy of self-distillation. Inspired by DINO and BYOL, we adopt a student-teacher architecture built upon the MAE encoder.
@@ 3.1 Datasets
To better evaluate our model, we conduct a series of diagnostic, comparative, and ablation studies on datasets including ImageNet-100, PACS and DomainNet (Peng et al., 2019). And all the detailed settings will be showed in the Appendix.
@@ 3.3 Comparison with State-of-the-Art
HyGDL 76.17 54.0 59.51 37.82 56.88 50.40 22.81 11.54 28.25
+: Use Imagenet transfer learnning.
MAE w/bss(sty): Added the bss module on top of the standard MAE (replaced the Fourier transform in bss with the AdaIN method).
Analysis of Results. As shown in Table 2, our HyGDL framework demonstrates strong performance on both PACS (56.88%) and the DomainNet subset (28.25%).(1) Superiority over SSL Baselines: HyGDL significantly surpasses standard SSL methods like SimCLR V2 and MAE, confirming the effectiveness of our explicit disentanglement approach for OOD robustness. (2) Validation of Principled Design: It also substantially outperforms heuristic-based approaches (MAE w/bss and MAE w/sty), validating our hypothesis that a principled mechanism is superior. (3) Competitive SOTA Performance: Crucially, while methods like CycleMAE leverage large-scale ImageNet pre-training, our HyGDL is pre-trained from scratch on only the source domain. Despite this significant disadvantage, it still achieves highly competitive results, outperforming many specialized UDG methods like DARLING. This highlights the exceptional data efficiency and powerful generalization capability learned by our approach.
@@ 4 Conclusion
Our extensive experiments validated the effectiveness of this principled approach. HyGDL significantly outperforms standard SSL and sophisticated heuristic-based baselines on challenging domain generalization benchmarks, confirming its ability to learn robust, style-invariant representations. More than just a new model, HyGDL serves as a proof-of-concept for a new class of robust SSL frameworks that move beyond mitigating symptoms to fundamentally altering the learning mechanism. For future work, the principle of explicit geometric disentanglement could be extended to other modalities, such as audio or medical imaging.
@@ Appendix A Appendix
A.1 Large Language Models
The authors gratefully acknowledge Gemini, a large language model from Google, for its valuable assistance in refining the manuscript’s structure, polishing the language, and offering insightful suggestions throughout the writing process.
@@ A.3.1 Diagnostic Experiment
For the diagnostic experiment, we follow the work of Geirhos et al. (2019) to create the IN-100-sty dataset based on ImageNet-100. The style transfer is performed using the AdaIN method with a pre-trained VGG-19. All training and evaluation are conducted on 1 NVIDIA RTX 4090 GPU. And the trainng settings are shown in Table 4:
Table 4: the Settings of Diagnostic Experiment
@@
A.3.2 Comparison Experiment
Table 5: the Settings of Comparison Experiment
@@ A.4.1 Asymmetric Generalization and Source Domain Challenges in Generative Pre-training
The results, presented in Table 6 reveal an asymmetry in generalization performance. In Setting A, both our HyGDL and the baseline MAE failed to generalize effectively. Conversely, in Setting B, our HyGDL demonstrated a clear advantage, successfully transferring knowledge to the schematic domains.
@@
HyGDL 20.71 27.11 27.60 25.14 50.40 22.81 11.54 28.25
+: Use Imagenet transfer learnning.
MAE w/bss(sty): Added the bss module on top of the standard MAE (replaced the Fourier transform in bss with the Adain method).
This asymmetry points to a fundamental limitation of MIM when the source domain is visually impoverished. Our investigation suggests the “Quickdraw” domain is the primary cause of the failure in Setting A. Generative-based pre-training relies on reconstructing rich visual context from partial inputs. When the source data is dominated by sparse, schematic line drawings like “Quickdraw”, the model is deprived of the necessary low-level texture and contextual cues (as shown in Figure 3). It fails to learn a rich enough feature vocabulary, leading to a catastrophic collapse in performance when transferring to visually dense domains like “Real” or “Painting”.
@@ A.4.2 Ablation studies
In this section, we provide additional ablation studies to further analyze the key components and hyperparameter sensitivity of our HyGDL framework. All experiments are conducted on the PACS dataset.

### v3

@@ 2.2.2 Identifying the Content Direction via Self-Distillation
While a conventional contrastive approach using an InfoNCE loss (Chen et al., 2020a) might seem applicable, we found it suboptimal for the fine-grained semantic alignment required between our strong positive pairs $(A,A^{\prime})$. This led us to reframe the problem: the core task is not to distinguish the pair from negatives, but to ensure their representations are identical in the content dimension.This objective aligns perfectly with the philosophy of self-distillation. Inspired by DINO and BYOL, we adopt a student-teacher architecture built upon the MAE encoder. By forcing the student encoder to predict the teacher’s output from a stylized view, the model acts as an information bottleneck. It is incentivized to discard the variable, non-essential information (style) and preserve only the shared, invariant information (content).
@@ 2.3 Explicit Disentanglement via Orthogonality
Our approach to disentanglement is predicated on the manifold hypothesis, which posits that our encoder learns to map high-dimensional images onto a more structured, lower-dimensional feature manifold. Assuming local linearity within this learned manifold, we can employ a principled geometric decomposition. This allows us to move beyond implicit heuristics and analytically separate content from style.
@@
For simplicity and interpretability, we instantiate content as a single invariant direction. In Appendix, we further extend this decomposition to multi-dimensional subspaces and observe consistent trends. We also include sensitivity analyses of different styles.
@@
We present the core principle using a single content direction for clarity; in practice, our main experiments use a two-dimensional content subspace ($k=2$) which yields the best trade-off (see Appendix A.5, Table 11).
@@ 3.1 Datasets
To better evaluate our model, we conduct a series of diagnostic, comparative, and ablation studies on datasets including ImageNet-100, PACS and DomainNet (Peng et al., 2019). Appendix details the settings and training overhead: HyGDL requires 40% more GPU-hours and higher peak memory than MAE (Table 12), though its inference cost is identical as it only uses the student encoder.
@@ 3.3 Comparison with State-of-the-Art
HyGDL($k=2$) 77.31 55.37 59.13 40.72 58.13 51.40 23.81 12.54 29.25
+: Use Imagenet transfer learning.
MAE w/bss and MAE w/sty: Added the bss module and the AdaIN style transfer module on top of the standard MAE, respectively.
Analysis of Results. As shown in Table 2, our HyGDL framework demonstrates strong performance on both PACS (58.13%) and the DomainNet subset (29.25%).(1) Superiority over SSL Baselines: HyGDL significantly surpasses standard SSL methods like SimCLR V2 and MAE, confirming the effectiveness of our explicit disentanglement approach for OOD robustness. (2) Validation of Principled Design: It also substantially outperforms heuristic-based approaches (MAE w/bss and MAE w/sty), validating our hypothesis that a principled mechanism is superior. (3) Competitive SOTA Performance: Crucially, while methods like CycleMAE leverage large-scale ImageNet pre-training, our HyGDL is pre-trained from scratch on only the source domain. Despite this significant disadvantage, it still achieves highly competitive results, outperforming many specialized UDG methods like DARLING. This highlights the exceptional data efficiency and powerful generalization capability learned by our approach.
@@ 4 Conclusion
Our experiments demonstrate that HyGDL effectively mitigates shortcut learning, outperforming standard SSL and heuristic baselines on multiple UDG benchmarks. Beyond proposing a model, this work provides a proof-of-concept for a principled framework based on explicit geometric disentanglement. Nonetheless, certain limitations remain: performance degrades in highly sparse domains such as Quickdraw, and results are sensitive to the choice of style-source, richer/style-diverse sources (e.g., Art) provide a stronger training signal (see Appendix A.4.1 and A.4.3). Future work may extend this principle to other modalities, including audio and medical imaging, and explore strategies to reduce sensitivity to domain characteristics. We hope this study encourages further research toward robust, interpretable, and generalizable representation learning.
5 Reproducibility Statement
To ensure the reproducibility of our findings, we provide the complete source code, pre-training scripts, and fine-tuning configurations in an anonymous Git repository here: https://anonymous.4open.science/r/HyGDL-45E1. All experimental details, including dataset processing, hyperparameters, and the specific three-stage training curriculum, are thoroughly documented in Section 3 and the Appendix.
@@ Appendix A Appendix
A.1 A Statement on Large Language Models Usage
A Large Language Model (LLM) was utilized as an assistant for manuscript preparation. Its contributions include refining the language and structure of the paper, as well as providing suggestions for articulating the theoretical underpinnings of our method. All core research ideas and experiments were conducted by the human authors.
@@ A.3.1 Diagnostic Experiment
For the diagnostic experiment, we follow the work of Geirhos et al. (2019) to create the IN-100-sty dataset based on ImageNet-100. The style transfer is performed using the AdaIN method with a pre-trained VGG-19. All training and evaluation are conducted on 1 NVIDIA RTX 4090 GPU. And the training settings are shown in Table 4:
Table 4: The settings of diagnostic experiment.
@@
Table 5: The settings of comparison experiment.
@@ A.4.1 Asymmetric Generalization and Source Domain Challenges in Generative Pre-training
The results, presented in Table 6 reveal an asymmetry in generalization performance. In Setting A, both our HyGDL and the baseline MAE failed to generalize effectively. Conversely, in Setting B, our HyGDL demonstrated a clear advantage, successfully transferring knowledge to the schematic domains. This asymmetry points to a fundamental limitation of MIM when the source domain is visually impoverished. Our investigation suggests the “Quickdraw” domain is the primary cause of the failure in Setting A. Generative-based pre-training relies on reconstructing rich visual context from partial inputs. When the source data is dominated by sparse, schematic line drawings like “Quickdraw”, the model is deprived of the necessary low-level texture and contextual cues (as shown in Figure 3). It fails to learn a rich enough feature vocabulary, leading to a catastrophic collapse in performance when transferring to visually dense domains like “Real” or “Painting”.
@@
HyGDL($k=2$) 21.71 28.11 28.60 26.14 51.40 23.81 12.54 29.25
+: Use Imagenet transfer learning.
MAE w/bss and MAE w/sty: Added the bss module and the AdaIN style transfer module on top of the standard MAE, respectively.
@@ A.4.2 Ablation studies
In this section, we provide additional ablation studies to further analyze the key components and hyperparameter sensitivity of our HyGDL framework. For these studies, we use our base model configuration with a content subspace dimensionality of $k=1$ as the benchmark, and all experiments are conducted on the PACS dataset.
@@
A.4.3 Sensitivity to Style Datasets
To evaluate the sensitivity of our framework to the source of stylization, we conducted experiments where the style images for AdaIN were sampled from different domains within the PACS dataset. The out-of-domain generalization performance is reported in Table 10. Note that we intentionally exclude the “Sketch” domain as a style source. As discussed previously (Appendix A.4.1), its sparse, line-based nature lacks the rich textural and color statistics required for effective stylization via AdaIN, which could introduce confounding variables into our sensitivity analysis.
Table 10: Out-of-domain accuracy on PACS when using different domains as the style source during pre-training. Best methods are highlighted in bold, and all reported accuracies are in percent (%).
PACS
Target Domain
Style Datasets photo art cartoon sketch avg.
photo 62.51 45.94 56.19 36.96 50.40
cartoon 57.54 47.85 55.85 40.49 50.43
art 76.17 54.0 59.51 37.82 56.88
Analysis. Our results, presented in Table 10, reveal an insightful dependency on the nature of the style source domain. Using the “Art” domain, which is characterized by diverse and rich textural patterns, as the style source yields the strongest OOD performance. This confirms that our framework is robust when provided with a sufficiently expressive stylization signal.
Conversely, using the “Cartoon” or “Photo” domains leads to a noticeable degradation in performance. We hypothesize two primary reasons for this. First, for domains like “Cartoon” with large, uniform color blocks, the feature statistics captured by AdaIN are less expressive, leading to a less effective stylization and thus a weaker training signal. Second, for the “Photo” domain, the stylistic difference between the source content image and a photo-stylized version is minimal. This makes the self-distillation task too simple, reducing the incentive for the model to learn a sharp and robust disentanglement of content and style.
This finding highlights a key insight: the effectiveness of our pre-training principle relies not just on style variation, but on variation that is sufficiently rich and distinct from the content source to drive robust disentanglement. This suggests that curating a diverse and stylistically expressive style dataset is an important factor for optimal performance.
A.5 Analysis of Content Subspace Dimensionality
To address the potential limitation of modeling content as a single direction, we extended our framework to support a multi-dimensional content subspace and investigated the effect of its dimensionality, $k$. Instead of using the average of student and teacher representations to define a single vector, we perform Singular Value Decomposition (SVD) on their cross-covariance matrix. The top-k left singular vectors are then used as an orthonormal basis for the k-dimensional content subspace. The content representation is subsequently computed by projecting the feature vector onto this subspace. And the process of identifying the k-dimensional content subspace and projecting onto it is formalized as follows.
Table 11: Out-of-domain accuracy on PACS with varying content subspace dimensionality ($k$). Best methods are highlighted in bold, and all reported accuracies are in percent (%).
PACS
Target Domain
Dimensionality k photo art cartoon sketch avg.
$k=1$ 76.17 54.0 59.51 37.82 56.88
$k=2$ 77.31 55.37 59.13 40.72 58.13
$k=4$ 75.69 52.83 59.01 38.10 56.41
Let $F_{A},F_{B}\in R^{B\times C}$ be the matrices of L2-normalized $[CLS]$ token representations for the teacher and student batches, respectively, where $B$ is the batch size and $C$ is the feature dimensionality. First, we compute the cross-covariance matrix between the two sets of representations:
$\sum_{AB}=\frac{1}{B}F^{T}_{A}F_{B}$
Next, we perform SVD on this matrix to find the principal directions of correlation:
$U,S,V^{T}=SVD(\sum_{AB})$
Where $U\in R^{C\times C}$ contains the left singular vectors as its columns.
The orthonormal basis for our k-dimensional content subspace, $v_{c}$, is formed by selecting the top-k left singular vectors from $U$:
$v_{c}=U_{:,:k}\in R^{C\times k}$
Finally, for any given feature vector $z\in R^{C}$, its content component $c\in R^{C}$ is computed by projecting $z$ onto this subspace:
$c=v_{c}v_{c}^{T}z$
The corresponding style component s is then derived as the orthogonal residual, s=zcs=z-c.
Analysis. According to the results on the PACS dataset are presented in Table 11, we found that increasing the dimensionality to $k=2$ yielded a notable improvement over the single-vector baseline ($k=1$). This suggests that a 2D subspace can capture a richer, more expressive representation of content than a single direction. However, further increasing the dimensionality to $k=4$ resulted in a performance degradation, falling below even the $k=1$ baseline. This finding indicates a ”sweet spot” for the content subspace’s capacity. While a single direction ($k=1$) might be too restrictive, a higher-dimensional space ($k=4$) may begin to capture redundant or even style-entangled features, effectively introducing noise into the content representation and harming the disentanglement process. This highlights the importance of choosing an optimal dimensionality and supports our choice of $k=2$ as the best configuration for this task.
A.6 Computational Cost Analysis
To quantify the computational overhead of our proposed framework, we compare its training efficiency with the standard MAE baseline. Both models were pre-trained on the PACS dataset (using 3 domains as source) for 1000 epochs on two NVIDIA RTX 4090 GPUs with a total batch size of 128. The results, calculated directly from the training logs, are summarized in Table 12.
Table 12: The comparison of computational cost, the primary overhead in HyGDL comes from the additional forward pass of the teacher encoder, cross-reconstruction and the SVD computation.
Model Hours GPU-Hours samples/sec GB perGPU
MAE 4.17 8.34 263.6 4.35
HyGDL 5.8 11.6 174 18.3
As shown in Table 12, our HyGDL framework introduces a notable computational overhead compared to the standard MAE. The total training time increases by approximately 40%, with a corresponding decrease in throughput. The peak GPU memory usage is also significantly higher. This overhead is an expected consequence of our training strategy, which involves multiple additional computations per step: a full forward pass through the teacher encoder for distillation, a second decoder pass for the cross-reconstruction objective, style-transfer based on AdaIN and the SVD computation for subspace identification.
It is crucial to note that this is a one-time pre-training cost. For all downstream applications and inference tasks, only the trained student encoder is utilized. Therefore, the inference cost of a HyGDL, in terms of latency, throughput, and memory footprint, is identical to that of a standard MAE model of the same architecture. We believe this investment in a more sophisticated pre-training process is a reasonable trade-off for the significant gains in out-of-domain generalization and the ability to learn robust, disentangled representations.